OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. y/3+y/4=5/6

OpenStudy (anonymous):

12

OpenStudy (anonymous):

4y/12+3y/12=10/12

OpenStudy (anonymous):

7y/12=10/12

OpenStudy (anonymous):

y=10/7

OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. 6x/11+9/2=3

OpenStudy (anonymous):

lcd is 22

OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation.x/2+4/5=9/12